OpenStudy (anonymous):

A sample of gas occupies 17 mL at -112 degreesC. What volume does the sample occupy at 70 degrees C?

OpenStudy (aaronq):

\(\dfrac{V_1}{T_1}=\dfrac{V_2}{T_2}\)

OpenStudy (aaronq):

hm i don't think the equations are working.. V1/T1=V2/T2 V=volume, T= temp (in Kelvin)